how many liters of 50% acid solution must be mixed with a 30% acid solution to get 180 liters of a 40% acid solution?
Answer by mananth(16949)

percent ---------------- quantity
50.00% ---------------- x
30.00% ---------------- 180-x
40.00% ---------------- 180

50x+30( 180-x)= 40*180

50x+5400-30x=7200

20x=1800
/20

x=90 ml of 50% solution